# Competition between group interactions and nonlinearity in voter dynamics on hypergraphs

### v1

@@ Abstract
Figure 5: (a) The dependence on the nonlinearity strength $q$ of the exit time $\tau$ for different values of the group size $s$ for hypergraphs with $N=10^{4}$ nodes. The markers are means of $10^{4}$ independent MC simulations on annealed $s$-uniform hypergraphs, and the analytical curves are from numerical solutions of the recursion relation. When $s=N$, the exit time $\tau$ grows exponentially quickly with $q$, making it problematic to depict it along with other cases. In the inset, we show the result for $s=N$ in an extended $q$ range. (b) The heat map for $\tau(s,q,N)$ that we obtain from numerical solutions of the recursion relation with $N=10^{4}$. For a given $s$, the cell with the red border has the optimal $\tau$.
Optimality in the exit time $\tau$—To further examine the complex interplay between the nonlinearity and group effect on the exit time $\tau$, we investigate how varying nonlinearity strength $q$ affects the GVM dynamics for specified values of the group size $s$ and hypergraph size $N$. In the absence of the group constraint (i.e., $s=N$), the leading-order expression of the exit time $\tau$ is
@@
This expression also applies to the $q$-voter model on a complete dyadic graph. Because $\tau$ increases with $q$, a stronger nonlinearity decelerates consensus [see Fig. 5(a) and its inset]. By contrast, with the most-constraining groups (i.e., $s=3$), $\tau$ decreases monotonically with $q$. To leading order, $\tau(q;N,s=3)\sim\frac{2^{q}}{2^{q-1}-1}\ln N$. A stonger nonlinearity accelerates consensus [see Fig. 5(a), red].
@@
There is a nontrivial tradeoff between these extreme situations. As we can see in Fig. 5(a), for a given hyperedge size $s$, there is an optimal nonlinearity strength $q^{*}$ with the minimum exit time $\tau$. We systematically investigate the tradeoff for many values of $s$ and $q$ [see Fig. 5(b)]. These computations reveal the global landscape of GVM dynamics and the associated geography of optimality.
@@
We now explain why we observe optimality. From the inset of Fig. 5(a), we see that considering the opinions of exactly $2$ neighbors is the most efficient way to achieve consensus. Increasing $q$ in Eq. (10) with $s=N$ reduces the probability that neighbors have unanimous opinions, which decreases the value of the drift function $v(\rho)$ and decelerates the approach to consensus. As $q$ grows logarithmically with $N$ [i.e., $q\sim{\cal O}(\ln N)$], the logarithmic scaling of $\tau$ in Eq. (10) eventually becomes a linear scaling $\tau\sim{\cal O}(N)$, which is comparable to the diffusive scaling for $q=1$. Moreover, when $s$ is finite, the probability that a node consults the same neighbor twice (instead of consulting 2 different neighbors) for $q=2$ is $1/(s-1)$, which is nonnegligible for small $s$. In this situation, a node only consults the opinion of $1$ neighbor, so it again effectively follows diffusive dynamics. More generally, the probability of diffusive dynamics from consulting just $1$ neighbor increases with decreasing $q$. Therefore, there is a “sweet spot” $q^{*}$ that minimizes $\tau$ when $2<q^{*}<{\cal O}(\ln N)$ between the two diffusive-dynamics maxima. The case of $s=3$ is notable exception to this; the maximum number of different neighbors is $2$, so $\tau$ decreasing indefinitely (although slowly) as $q$ increases.
We can also explain the presence of optimality in the simplicial GVM in Fig. 2 as a competition between the diffusive dynamics from dyadic edges (which dominates as $\langle s\rangle\downarrow 2$ and $\alpha\to\infty$) and the small probability of unanimity in large hyperedges (which dominates as $\langle s\rangle\to\infty$ and $\alpha\downarrow 2$). Following the same logic as above, we again obtain an optimum in the opinion dynamics.
@@
Conclusions—We formulated and analyzed a group-driven voter model (GVM) that accounts for the effects of both polyadic and nonlinear interactions within groups. A larger nonlinearity strength $q$ leads to faster consensus in the GVM than in conventional VMs, which exhibit diffusive dynamics. This acceleration of consensus formation depends on the interplay between the nonlinearity strength $q$ and the group size $s$ of hypergraphs. Through mean-field calculations and Monte-Carlo simulations, we demonstrated that the exit time scales logarithmically with system size and that there is an optimum value $q^{*}$ of the nonlinearity strength $q$ that minimizes the exit time. This optimality emerges from a competition between diffusive dynamics when both $q$ and $s$ are small and a slow drift when both $q$ and $s$ are large. This emergent group effect cannot arise in dyadic networks, providing justification for the analysis of dynamics on polyadic networks.
We also apply our analytical approach to several variants of our GVM (see the SM SM ): the simplicial GVM, the GVM without allowing duplicate choices, and a GVM with edge-update dynamics in which we simultaneously update the opinions of all nodes that are attached to a hyperedge. In all of these cases, the exit time scales logarithmically with system size, illustrating the robustness of our main theoretical results SM .
@@ Acknowledgements.
Acknowledgments—We thank Sid Redner for useful conversations and suggestions. This work was supported in part by National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) [No. 2020R1A2C2003669 (K-IG) and No. 2020R1I1A3068803 (BM)]; by a KIAS Individual Grant [No. CG079902] at Korea Institute for Advanced Study (D-SL); by the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (MAP); and by Agencia Estatal de Investigación (AEI, MCI, Spain) MCIN/AEI/10.13039/501100011033 and Fondo Europeo de Desarrollo Regional (FEDER, UE) under Projects APASOS [PID2021-122256NB-C21] and the Maria de Maeztu Program for Units of Excellence in R&D grant [CEX2021-001164-M] (MSM).

### v2

@@ Abstract
Figure 5: (a) The dependence on the nonlinearity strength $q$ of the exit time $\tau$ for the GVM for different values of the group size $s$ for hypergraphs with $N=10^{4}$ nodes. The markers are means of $10^{4}$ independent MC simulations of the GVM on annealed $s$-uniform hypergraphs, and the analytical curves are from numerical solutions of the recursion relation (S21) in the SM SM . When $s=N$, the exit time $\tau$ grows exponentially quickly with $q$, making it problematic to depict it along with other cases. In the inset, we show the curve for $s=N$ in an extended $q$ range. (b) The heat map for $\tau(s,q,N)$ that we obtain from numerical solutions of the recursion relation (S21). For a given $s$, the cell with the red border has the optimal $\tau$.
Optimality in the exit time $\tau$—To further examine the interplay between the nonlinearity and group effects on the exit time $\tau$, we investigate how varying the nonlinearity strength $q$ affects the GVM dynamics for specified values of the group size $s$ and hypergraph size $N$. In the absence of the group constraint (i.e., $s=N$), the leading-order expression of the exit time $\tau$ is
@@
This expression also applies to the $q$-voter model on a complete dyadic graph. Because $\tau$ increases with $q$, a stronger nonlinearity decelerates consensus. [See Fig. 5(a) and its inset.] By contrast, with the most-constraining groups (i.e., $s=3$), the exit time $\tau$ decreases monotonically with $q$. To leading order, $\tau(q;N,s=3)\sim\frac{2^{q}}{2^{q-1}-1}\ln N$. A stonger nonlinearity accelerates consensus. [See the red curve in Fig. 5(a).]
@@
There is a nontrivial tradeoff between these extreme situations. As we can see in Fig. 5(a), for a given hyperedge size $s$, there is an optimal nonlinearity strength $q^{*}$ with the minimum exit time $\tau$. We systematically investigate the tradeoff for many values of $s$ and $q$ [see Fig. 5(b)]. These computations reveal the global landscape and optimality of GVM dynamics.
@@
We now explain why we observe optimality. From the inset of Fig. 5(a), we see that considering the opinions of exactly $2$ neighbors is the most efficient way to achieve consensus. Increasing $q$ in Eq. (10) with $s=N$ reduces the probability that neighbors have unanimous opinions, which in turn decreases the drift function $v(\rho)$ and decelerates the approach to consensus. Because $q$ grows logarithmically with $N$ [i.e., $q\sim{\cal O}(\ln N)$], the logarithmic scaling of $\tau$ in Eq. (10) eventually becomes a linear scaling $\tau\sim{\cal O}(N)$, which is comparable to the diffusive scaling for $q=1$. However, the probability that a node consults the same neighbor twice (instead of consulting 2 different neighbors) for $q=2$ is $1/(s-1)$, which is not negligible for small $s$. In this situation, a node only consults the opinion of $1$ neighbor, so it again effectively follows diffusive dynamics. More generally, the probability of diffusive dynamics from consulting just $1$ neighbor increases with decreasing $q$. Therefore, there is a “sweet spot” $q^{*}$ that minimizes $\tau$ between the two diffusive-dynamics maxima. That is, $2<q^{*}<{\cal O}(\ln N)$. The case $s=3$ is a notable exception. When $s=3$, the maximum number of different neighbors is $2$, so $\tau$ decreases indefinitely (although slowly) as $q$ increases.
When $q$ is fixed, increasing $s$ towards $N$ reduces the probability of unanimity, as the number of distinct neighbors that are chosen increases, and decreasing $s$ towards $2$ increases the probability of consulting just $1$ neighbor. Both situations lead to an increase in the exit time $\tau$. Therefore, there exists an optimal $\tau$. An equivalent explanation of the presence of optimality in the simplicial GVM (with $q=s-1$ and duplicate choices disallowed) in Fig. 2 is as a competition between diffusive dynamics from dyadic edges (which dominates as $\langle s\rangle\downarrow 2$ and $\alpha\to\infty$) and the small probability of unanimity in large hyperedges (which dominates as $\langle s\rangle\to\infty$ and $\alpha\downarrow 2$). The exit time $\tau$ increases as one approaches either of these limits, so there is an optimal $\tau$.
@@
Conclusions—We formulated and analyzed a group-driven voter model (GVM) that accounts for the effects of both polyadic interactions and nonlinear interactions within groups. A larger nonlinearity strength $q$ leads to faster consensus in the GVM than in conventional VMs, which exhibit diffusive dynamics. This acceleration of consensus formation depends on the interplay between the nonlinearity strength $q$ and the group size $s$ of hypergraphs. Through mean-field calculations and Monte-Carlo simulations, we demonstrated that the exit time scales logarithmically with system size and that there is an optimal value $q^{*}$ of the nonlinearity strength $q$ that minimizes the exit time. This optimality emerges from a competition between diffusive dynamics when both $q$ and $s$ are small and a slow drift when both $q$ and $s$ are large. This emergent group effect cannot arise in dyadic networks.
We also apply our analytical approach to several variants of our GVM (see the SM SM ): a simplicial GVM, a GVM without allowing duplicate choices, and a GVM with edge-update dynamics in which we simultaneously update the opinions of all nodes that are attached to a hyperedge. In all of these cases, the exit time scales logarithmically with system size, illustrating the robustness of our main theoretical results SM . To further examine the robustness of our results, it is also important to consider additional phenomena. For example, our analysis did not account for heterogeneities in the degree distribution (where the degree of a node is the number of hyperedges it is in). Our preliminary Monte-Carlo calculations SM illustrate that the degree distribution can influence the exit time. Therefore, it will be useful to generalize our mean-field framework to a degree-based mean-field theory Sood and Redner (2005); Sood et al. (2008); Vazquez and Eguíluz (2008) to study their effects analytically. It will also be useful to extend our GVM to encompass more realistic aspects (see, e.g., Redner (2019)) of opinion dynamics. There have been studies of optimal group and team sizes in social psychology Bond (2005); Wheelan (2009), and further studies of our GVM and its generalizations may yield interesting insights about these phenomena.
@@ Acknowledgements.
Acknowledgments—We thank Sid Redner for useful conversations and suggestions. This work was supported in part by National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) [No. 2020R1A2C2003669 (K-IG) and No. 2020R1I1A3068803 (BM)], by a KIAS Individual Grant [No. CG079902] at Korea Institute for Advanced Study (D-SL), by the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (MAP), and by Agencia Estatal de Investigación (AEI, MCI, Spain) MCIN/AEI/10.13039/501100011033 and Fondo Europeo de Desarrollo Regional (FEDER, UE) under Projects APASOS [PID2021-122256NB-C21] and the Maria de Maeztu Program for Units of Excellence in R&D grant [CEX2021-001164-M] (MSM).